OpenStudy (anonymous):

Solve using separation of variables: f''(x)g(y) + (1/x)f'(x)g(y) + (1/x^2)f(x)g''(y)=0

OpenStudy (anonymous):

Divide by f(x)g(y)!

OpenStudy (anonymous):

Hahahahahah... I'm embarrassed via the internet right now but okay thanks

OpenStudy (anonymous):

don't try too hard... you might break a sweat :)

OpenStudy (anonymous):

and multiply by x^2

OpenStudy (anonymous):

try to get this \[\large \frac{x^2 f''(x)+xf'(x)}{f(x)}=-\frac{g''(y)}{g(y)}=\lambda\]

OpenStudy (anonymous):